The school that Jessica goes to is selling tickets to the annual dance competition. On the first day of ticket sales the school sold 9 senior citizen tickets and 14 child tickets for a total of $227. The school took in $171 on the second day by selling 3 senior citizen tickets and 12 child tickets. What is the price each of one senior citizen ticket and one child ticket?
Solve by elimination.

Answer by Boreal(15235)

9S+14C=227
3S+12C=171
-9S-36C=-513, multiplying the second by -3 to eliminate S
-22C=-286
C=$13 per child
S=$5 per senior, substituting into either equation